Status Report

NASA Space Station On-Orbit Status 15 September 2004

By SpaceRef Editor
September 15, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 15 September 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

Update on Elektron:  The Elektron O2 generator is still Off.  Further troubleshooting by the crew was put on hold today to give ground experts time for planning the next steps. [Yesterday, the crew successfully purged both water and air through the H2 and O2 lines.  They were then advised by TsUP/Moscow to disassemble the purging equipment and reassemble the H2 and O2 lines in the same configuration as before, i.e., without the gas analyzer on the H2 line.  Specialists are now assessing whether or not the next steps should include replacing the liquid unit BZh-7 (blok zhidkostnoi 7) with the older BZh-5.  A decision is expected either tonight or tomorrow morning.]


Service Module Atmosphere Revitalization Subsystem, Book 2, Mission Operations Directorate, 9 October 2000 [Acrobat] According to this document’s introduction “This book contains information for the crew about procedures and rules for the atmosphere revitalization subsystem, Elektron, Vozdukh, Micropurification Unit, and Fire Detection and Suppression Subsystem operations, as well as their schematic and operation logic.”

Update on cabin atmosphere:  ISS cabin pressure was repressurized with 13 mmHg of fresh oxygen from Progress 15P storage at ~11:00am EDT.  ppO2 rose by 12.2 mmHg. [About seven hours before the repress, the U.S. MCA (major constituents analyzer) was activated, and a “zero” calibration was performed on it (9:35am) to ensure valid O2 readings for the repress.  Later in the day, the MCA was returned to LEM (life extending mode) to preserve its ion pump.]

CDR Gennady Padalka prepared the Soyuz-214 communications system for the third and last day of VHF2 checkout with U.S. ground sites.  The testing was successfully conducted with White Sands Complex (WSC) at 4:40am. [The location of the Soyuz VHF2 uplink & downlink frequencies in the VHF aeronautical comm band restrict NASA VHF ground stations operations on Soyuz VHF2 to spacecraft emergencies only.  Due to the emergency nature of any possible Soyuz VHF2 support from NASA ground sites during Soyuz solo-flight phases, NASA has obtained permission from the FAA and other regulatory agencies to conduct a one-time test with each NASA VHF ground station.  This was a significant milestone for the ISS Program and for the NASA Network.  It required close coordination with the NASA HQ Spectrum officials, NTIA, FAA, MCC, ISS program, Flight crew and Network personnel.]

Later, Padalka began the long-awaited work on improving the acoustic noise level in the Russian segment (RS), which will replace hard air ducts and hard ventilator mounts with soft ducts and isolated mountings.  First part of this activity was deferred yesterday.  Before and after the muffler installation, Padalka was to take sound level readings with the Russian “Shumomer” noise measurement instrument, including background noise without fans running.  The noise data were later transferred to the MEC (medical equipment computer, see also below) for downlink. [The newly delivered cylindrical disk-shaped vibration isolator cartridges are designed for installation on five fans of the Service Module (SM) ventilation system, located behind wall panels, viz. panel 307 (VPO5 and VPO6), 129 (VPO10), 131 (VPO11), 130 (VPO12) and 126 (VSZP-1).]

In the Lab module, FE/SO Mike Fincke powered up the HRF GASMAP (Human Research Facility/Gas Analyzer System for Metabolic Analysis Physiology) and its laptop for the regular routine 30-day health check.  Afterwards, the equipment was turned off again. [The run was shortened by five hours because of maintenance work being performed at JSC on the Telemetry Server (which impacted the HRF team’s ability to see the GASMAP telemetry downlink), but it sufficed to satisfy minimum health check objectives.]

In the SM, the CDR deactivated the IK0501 gas analyzer (GA) and exchanged its BF carbon dioxide (CO2) filter assembly with a new unit from FGB stowage (replaced last: 8/5).  GA was reactivated and the spent BF stowed for return to Earth on Soyuz.  [IK0501 is an automated system for measuring CO2, O2, and H2O in the air as well as the flow rate of the gas being analyzed.]

Going by the uplinked recovery plan for the failed MEC (medical equipment computer) chassis, Mike Fincke removed the resident hard disk drive (HDD) from the MPSD2 (Multi-Purpose Support Disc-2) laptop and replaced it with the MEC hard drive.  The now empty laptop shell (#1001) was stored along with the removed HDD.  The IP (Internet Protocol) phone function of the MPSD was re-established on another laptop (SSC5).

Using the new MEC, Mike took his fifth periodic On-Orbit Hearing Assessment (O-OHA) test, a NASA environmental health systems examination to assess the efficacy of acoustic countermeasures. [The O-OHA audiogram test involves minimum audibility measurements for each ear over a wide range of frequencies and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter).  To conduct the testing, the experimenter is supported by special EarQ software on the MEC (medical equipment computer).  The baseline test is required not later than about Flight Day 14 for each new Expedition and is then performed once per month.]

The FE conducted the periodic verification of IMV (intermodular ventilation) airflow between the U.S. (USOS) and Russian segment (RS) using the Velocicalc meter.  He also took CO2 readings with the CDMK (CO2 monitoring kit). [There is no direct measurement of airflow except as reflected by differences in atmosphere partial pressures measured between the RS and USOS.  ppCO2 (CO2 partial pressure) is a good yardstick since an increasing ppCO2 in the Lab not reflected in the SM indicates that Vozdukh is not receiving the air from the Lab at an efficient rate.  Periodic air flow degradation checks support establishing a most effective fan cleaning schedule.]
Fincke also performed a pressure check on the PMA-2 (pressurized mating adapter #2) with a scopemeter, after attaching the ISA (internal sample adapter) equipment to the MPEV (manual pressure equalization valve) in the Lab forward hatch.  Later, the setup was removed and temporarily stowed for similarly use on PMA-3 and also for the Lab nadir window activities scheduled for Friday (9/17).

Afterwards, Mike completed the routine inspection of the Lab starboard heat exchanger of the CCAA (common cabin air assembly), checking for water that may have accumulated on the heat exchanger, one of the local “cold spots” that could attract condensate.

Gennady conducted the periodic inspection of the BIO-5 Rasteniya-2 (“Plants-2”) experiment which researches growth and development of plants under spaceflight conditions in the Lada-5 greenhouse.

Mike did the routine maintenance of the SM’s SOZh environment control & life support system, today including the weekly inspection of the BRPK air/liquid condensate separator apparatus.  He also prepared the regular IMS (inventory management system) delta file for export/import to the IMS databases and completed the routine status checkup of the autonomous PCG-STES010 (Protein Crystal Growth-Single Locker Thermal Enclosure System) payload in the Lab (done every Monday, Wednesday and Friday).

Both crewmembers completed their daily 2.5-hr. aerobic/anaerobic workout program on TVIS treadmill, CEVIS bike, RED exerciser and VELO cycle with bungee cord load trainer.

Later in the afternoon, the CDR set up the equipment for the Russian MedOps biochemical blood analysis MO-11, consisting of the Reflotron-4 analyzer, with accessories, power supply and Reflotron-4 kit.  After setup, the instrument was deactivated and left fully configured at the work site until tomorrow morning. [Gennady will undergo the 2.5-hr. tests tomorrow (8/20), preceded by imbibing 250 ml of warm water or plain (unsugared) tea 20 minutes before taking the blood samples.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

The crew was thanked for yesterday’s successful ADUM (Advanced Diagnostic Ultrasound in Micro-G) session, which resulted in another set of cardiac (heart) and thoracic (chest) ultrasound scans on Mike while wearing ECG electrodes.

After ISS attitude control was handed over to the RS motion control system at 9:35am, the planned uplink of GNC (guidance, navigation & control) software version R4 for the GNC MDMs (multiplexer/demultiplexers, computers) took place at 10:00am. [Today, the first day of uplink, the backup GNC (GNC-1) was transitioned to diagnostics mode and loaded with the vers.R4 software.  The primary GNC-2 was then swapped to backup, while the newly loaded MDM became Primary.  MCC-Houston then re-initialized and configured the new Backup (GNC-2), keeping the R3 software on it for a 24-hour wait.  Tomorrow, GNC-2 will be transitioned to diagnostics mode and loaded with R4.  The end result of the two-day activity will be GNC-1 as primary and GNC-2 as backup GNC MDM.  The vers.R4 software (as well as the current R3 load) are known to contain an error that, in an extremely unlikely event, could result in an inaccurate GPS attitude solution.]

Prior to the control handover, the U.S. P6 solar array wings (SAWs) were set to “Directed” position mode, with the BGAs (beta gimbal assemblies feathered at 135 deg (4B) and 225 deg (2B).  Later, at 2:00pm, they were returned to “Autotrack”. [While in XPOP (x-axis perpendicular to orbit plane), the P6 BGAs are nominally configured into Autotrack mode, in which the SAWs automatically track the sun.  However, during the sunrise and sunset portions of each orbit, the SAWs’ surface area produces a large amount of drag.  To reduce this drag and thereby save propellant, a “bias” angle is used to keep the arrays as “edge on” to the velocity vector as possible while still producing the power required for nominal operations.  Currently this bias angle is set to 47 degrees, and will remain in place until the attitude change to LVLH (local vertical/local horizontal) on 9/17.]

Overnight the ground uplinked an updated list of 19 “yellow tag” items, including their constraints on usage aboard ISS, for an on-orbit review by the crew.  Before a yellow-tagged item can be used, the ground needs to be consulted. [Yellow tags, more formally called “uncertified dual ops tags”, are used to identify items not certified for ISS Operations (certification and/or paperwork not complete prior to launch); items which have IP (international partner) segment-specific certification (can be used in one IP segment but should not be used in anther IP segment); items that could pose a safety hazard; and items that are broken or expired.  Blank yellow tags are flown so hardware can be tagged on-orbit as necessary.]

The total number of CEO (crew earth observations) images for Increment 9 as of last Friday, 9/10, at COB, is 15,072. 
Expedition 9 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Today’s CEO photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in ram), were Internal waves, Patagonian Shelf (clear weather windows over southern South America provide an opportunity for internal wave photography.  The sunglint point was on the left and slightly ahead of track), Salamat Basin fans, Chad (large-scale oblique images of this fan will be useful to augment existing detailed photography. The fan was to the left of track during the ISS overpass), and Smoke plume, central South America (Dynamic Event.  Review of weather satellite data indicates some interesting plume patterns due to local weather systems.  Looking to the right of track for a thinned neck of the main plume interacting with a cloud system. Inclusion of recognizable land features (i.e. the Andes) will help speed the cataloging of images).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

Major upcoming events:

  • Reboost — 9/22? (phase angle correction for 9S)
  • Soyuz 9S launch — TBD (10/14?) (w/Leroy Chiao, Salizhan Sharipov, Yuri Shargin);
  • Soyuz 9S dock — TBD (10/16?);
  • Soyuz 8S undock — TBD (10/23?) (w/Gennady Padalka, Michael Fincke, Yuri Shargin);
  • Soyuz 8S land — TBD (10/23?);
  • Soyuz 9S port relocate — 11/18;
  • Progress 15P undock – 11/23;
  • Progress 16P launch — 11/24;
  • EVA-12 – 12/28;
  • Progress 16P undock – 1/29/05;
  • Progress 17P launch – 1/30/05;
  • EVA-13 – 2/21/05;
  • Shuttle/LF1 launch – NET 3/6/05;
  • Shuttle/LF1 undock – NET 3/16/05.

U.S. & Russian Segment Status  (as of today, 1:37pm EDT)
Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is Off.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is On.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is Off.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off.  SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; new replaceable condensate removal line installed on 9/9).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 750; temperature (deg C) — 25.7; ppO2 (mmHg) — 159.6; ppCO2 (mmHg) — 3.8.
  • SM Transfer Compartment:  Pressure (mmHg) — 741; temperature (deg C) — 20.1.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 22.3.
  • Node:  Pressure (mmHg) — 740.8; temperature (deg C) — 24.3 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 747.7; temperature (deg C) — 24.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 747.9; temperature (deg C) — 26.7; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, “sun slicer”, i.e., drag reduction-biased by 47 deg angle on 2B & 4B).
  • SM batteries:  Battery #6 is off; all other batteries (7) are on line in “Partial Charge” mode.
  • FGB batteries:  Battery #3 is off line; battery #4 is in “Cycle” mode; all other batteries (4) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems (C&DH)

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-1 MDM (R4) is prime; GNC-2 (R3) is backup (see R4 upgrade, above).
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System (PS):

  • Total propellant load available: 4408 kg (9718 lb) as of 9/8/04;  [SM(552) + FGB(3388) + Progress M(468)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems (ACS):

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 180.5 deg, pitch: -6.9 deg., roll: 0 deg]), with CMG TA (thruster assist) until 9/17.

Communications & Tracking Systems (C&T):

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 9:26am EDT [= epoch]):

  • Mean altitude — 359.7 km
  • Apogee height — 363.4 km
  • Perigee height — 356.1 km
  • Period — 91.74 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0005426
  • Solar Beta Angle — 40.6 deg (magnitude decreasing)
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 150 m
  • Revolutions since FGB/Zarya launch (Nov. 98)  — 33251


ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.